OpenStudy (anonymous):

Please simplify the following equation:

OpenStudy (gorv):

i think it should be 12*x^(1/2)

OpenStudy (gorv):

@bobobobobb plzzzz check the Q

OpenStudy (anonymous):

\[4x ^{3/2} - 12x^{1/2} + 8x ^{-1/2}\]

OpenStudy (anonymous):

Sorry its supposed to be this.^

OpenStudy (gorv):

ok multiply equation by x^1/2

OpenStudy (gorv):

4x^3/2 * x^1/2 -12x^1/2 * x^1/2 + 8 X^(-1/2) * x^1/2 =4x^2-12x+8 4x^2-8x-4x+8=4x *(x-2)-4(x-2)=(x-2)*(4x-4)= 4*(x-1)*(x-2)

OpenStudy (gorv):

@bobobobobb got it buddy???